• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name B. Craver

    Log entry time 16:05:50 on October19,2006

    Entry number 187666

    keyword=day shift summary

    08:00 Problems with left and right arm NMR
    08:30 Beam is down for the beam studies. We request Controlled Access. Ed and crew enter the hall. They will look at the spectrometer. We run the DAQ on cosmics during the access.
    12:30 The techs finish working and leave the hall.
    12:55 Hall A is back to Power Permit



    Run Summary:

    Run No. Type Ev's
    1529 Cosmics 200k
    1530 Cosmics 200k
    1531 Cosmics 200k

    15:11 Beam Production
    1532 B4 2M
    1533 B4 2M
    1534 B4 2M

    Continuing to next run